OpenStudy (anonymous):

hello I can't success to solve this 2^(1/2)cos(x+(pi/6))>=-1 I use parametric form but I can't success how i must do?

OpenStudy (anonymous):

\[2^{1/2}\cos (x+\pi/6)\ge-1\] is this the inequation??

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

well, do u need general solution or solution between -pi to pi???

OpenStudy (anonymous):

i tried with that site but it give me only the solution I need to know how to solve this inequation

OpenStudy (anonymous):

oo.. sry i'm not good in trigonometry inequalities

OpenStudy (anonymous):

xd this is the same problem for me :p do you know somebody that could help me with this?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=h--b_b5gW_0 try to watch...

OpenStudy (anonymous):

actually if what i want to know step by step i search youtube...

OpenStudy (anonymous):

I know how solve it but the problem it's that doesn't work

OpenStudy (anonymous):

solution are : ]-pi/12+2kpi, 7pi/12+2kpi[

OpenStudy (anonymous):

problem is how i can get this

OpenStudy (anonymous):

.. sry.. i dont know who else can solve.. but i know 1 person... isn't online now.. satelite.... 100 lvl in math.. always online in 8-11pm (GMT+8 malaysia time zone)

OpenStudy (anonymous):

sorry, i gave the wrong answer at first

OpenStudy (anonymous):

thanks for your advice neoh147 arnab09 you can't help me?

OpenStudy (anonymous):

i can.. i guess... but gotta take more time :(

OpenStudy (anonymous):

ok so advice me only how i can solve it for example say me if i must solve the cos addition? and then i must use parametric form?

OpenStudy (anonymous):

well, i am trying to solve drawing quadrants

OpenStudy (anonymous):